Our team identified an issue with the uplink on one of our switch.
Services impacted are located in room 3-2 racks D17 and D18
09/10/2018 1130Z (14 00 LT) - The problem is solved. A replacement of switch was necessary. Affected servers should be responding properly now.
09/10/2018 1130Z (13 30 LT) - Our engineers are already investigating, issue is being fixed. Next update at 14 LT.
We will keep you updated through this status.